OpenStudy (anonymous):

In the following reaction, how many grams of ammonia (NH3) will react with 27.8 grams of nitric oxide (NO)? 4NH3 + 6NO d 5N2 + 6H2O The molar mass of ammonia is 17.0337 grams and that of nitric oxide is 30.01 grams. A.23.7 grams B. 10.5 grams C. 73.5 grams D. 32.7 grams

OpenStudy (cuanchi):

NO= 14+16= 30 g/mol NH3 = 14+3=17 g/mol n= m/MM= 27.8 g/30 g/mol= 0.9266 moles of NO 0.9266 moles of NO x (4 mol NH3/6 mol NO) x (17 g NH3/mol/1mol NH3)= ????? g NH3